# Unleashing the Potential of Large Language Models for Text-to-Image Generation through Autoregressive Representation Alignment

### v2

@@ Abstract
We present Autoregressive Representation Alignment (ARRA), a new training framework that unlocks global-coherent text-to-image generation in autoregressive LLMs without architectural changes. Unlike prior work that requires complex architectural redesigns, ARRA aligns LLM hidden states with visual representations from external visual foundational models via a global visual alignment loss and a hybrid token, <HYBNEXT>. This token enforces dual constraints: local next-token prediction and global semantic distillation, enabling LLMs to implicitly learn spatial and contextual coherence while retaining their original autoregressive paradigm. Extensive experiments validate ARRA’s plug-and-play versatility. When training from text-generation-only LLMs or random initialization, ARRA reduces FID by 25.5% (MIMIC-CXR), 8.8% (DeepEyeNet), and 7.5% (ImageNet) for advanced autoregressive LLMs like Chameleon and LlamaGen, all without framework modifications. For domain adaption, ARRA aligns general-purpose LLMs with specialized models (e.g., BioMedCLIP), achieving an 18.6% FID reduction over direct fine-tuning on medical imaging (MIMIC-CXR). By demonstrating that training objective redesign—not just architectural innovation—can resolve cross-modal global coherence challenges, ARRA offers a complementary paradigm for advancing autoregressive models. The code is available at https://github.com/xiexing0916/ARRA.
Figure 1: ARRA enables high-quality text-to-image generation through redefined training objective. (a)(c) Traditional next-token prediction (NTP)-based LLMs rely solely on the autoregressive loss (AR loss) of the next token <NEXT> for local constraints. (b)(d) ARRA constructs a hybrid token <HYBNEXT>, which is aligned by introducing external global visual representation, ensuring that <HYBNEXT> is constrained both locally by the AR loss and globally by the global visual alignment loss (GVA loss). (e)(f) ARRA demonstrates advantages in semantic consistency and visual continuity.
@@ 1 Introduction
Large language models (LLMs) [1, 4, 34, 36, 37] and subsequent multimodal LLMs (MLLMs) [28, 24, 51, 55, 13, 29] has revolutionized the field of generative AI. These models, built on the autoregressive paradigm, demonstrate remarkable scalability and generalization capabilities in complex reasoning and understanding tasks through a simple yet powerful next-token prediction framework.
@@
Inspired by the success of LLMs, recent work has sought to replicate their autoregressive “next-token prediction” paradigm for text-to-image generation. Early works like DALL·E [41], Parti [60] and LlamaGen [48] attempted to directly adopt the next-token prediction paradigm of LLMs for image generation by treating images as sequences of discrete tokens. However, while the “next-token prediction” paradigm excels in text—where local dependencies naturally align with sequential structure—it struggles to bridge the significant cross-modal gap between language and images. As shown in Fig. 1(c), optimizing only for local next-token prediction forces the model to focus on isolated token-level features, neglecting the global coherence required for spatially structured visual content. This may result in images with 1) fragmented outputs—e.g., misaligned ribs in X-rays, where fine-grained details fail to harmonize into a unified whole, and 2) semantic mismatches, as shown in Fig. 1(e), where global information is not maintained, leading to inconsistencies in the generated visual content.
@@
Recognizing this limitation, recent efforts aim to inject global constraints into autoregressive frameworks to fully unlock the potential of LLMs in image generation [62, 58, 52]. For example, Transfusion [62] and Show-O [58] introduce bidirectional attention mechanisms to model global image structure through patch diffusion and mask token modeling, respectively. While these methods achieve remarkable results—generating high-quality images and demonstrating LLMs’ potential for unified multimodal tasks—they rely on architectural modifications such as cross-modal attention layers or grafted diffusion modules. Though effective, such adaptations often depart from standard LLM frameworks, limiting their compatibility with pretrained LLMs that excel under pure autoregressive paradigms. For instance, repurposing an off-the-shelf LLM for text-to-image generation would require retraining with these specialized components, forfeiting the benefits of existing scaling laws and generalization capabilities. This practical constraint raises a critical question: Can we unlock the full potential of LLMs for text-to-image generation without altering the original architecture or inference mechanism?
@@
We address this by proposing Autoregressive Representation Alignment (ARRA), a novel training framework that redefines how LLMs learn text-to-image generation. Unlike prior work that modifies architectures (e.g., adding attention layers or diffusion modules), ARRA preserves the original LLM framework while injecting global constraints directly into the training objective. Our key insight is simple: global coherence need not come from architectural complexity; it can instead emerge from a redefined training paradigm. Specifically, ARRA augments the standard autoregressive loss with a global visual alignment loss that aligns the LLM’s latent representations with semantic guidance from pretrained foundational models (Fig. 2). To bridge local and global learning, we introduce a hybrid token, <HYBNEXT>, which serves as a bidirectional anchor: Locally, it predicts the next token via standard codebook indices. Globally, its latent embedding aligns with compressed visual features extracted from external models (e.g., BioMedCLIP [61] or MedSAM [33]) via our novel global visual alignment loss. By distilling rich semantic features (e.g., spatial relationships, object coherence) from external models into the <HYBNEXT> token during training, ARRA enables autoregressive sequences to implicitly learn global structure. Crucially, this alignment occurs only during training, leaving the LLM’s inference process untouched and preserving its inference-time efficiency.
Our experiments validate ARRA’s versatility across diverse domains, including medical imaging and natural scene synthesis, where it achieves state-of-the-art Fréchet Inception Distance (FID) and CLIP-Score improvements without architectural changes. Notably, ARRA’s plug-and-play nature complements existing innovations: when paired with architectures like Chameleon [50] or Llamagen [48], it further enhances output quality, demonstrating additive benefits. Furthermore, ARRA demonstrates its capability for domain adaptation. By infusing knowledge from domain-specific foundational models (e.g., BioMedCLIP [61] or MedSAM [33]) into LLMs with image generation capabilities, such as Lumina-mGPT [27], its generation performance improves significantly compared to direct fine-tuning. This indicates that ARRA can facilitate the transition of models from a general-purpose generative framework to domain-specific generative models.
@@
(i) We introduce Autoregressive Representation Alignment, a novel training framework that redefines how LLMs learn text-to-image generation by decoupling global structure learning from model design. By aligning training objective with external representations, ARRA resolves local dependency limitations in LLMs while retaining original architectures and inference efficiency.
(ii) We introduce the <HYBNEXT> token, a novel mechanism that bridges local next-token prediction with global semantic alignment via distillation from external models (e.g., BioMedCLIP [61] or MedSAM [33]), enabling implicit learning of spatial and contextual relationships.
(iii) We provide a detailed experimental analysis, offering comprehensive insights into the selection of external representations, alignment tokens, and alignment depth. These findings serve as a practical guide for the effective utilization of representation alignment.
(iv) We validate ARRA’s ability to 1) enhance advanced autoregressive LLMs (e.g., Chameleon, LlamaGen) without framework modifications: the FID reduces by 25.5%, 8.8%, 7.5% on the MIMIC-CXR, DeepEyeNet and ImageNet datasets, and 2) seamlessly adapt general-purpose models to specialized domains (e.g., medical imaging) via domain knowledge infusion: the FID reduces by 18.6% on the MIMIC-CXR.
Figure 2: Proposed ARRA Framework. We define the next token predicted in the autoregressive sequence as the “HYBRID next token”, denoted as <HYBNEXT>. During training, <HYBNEXT> is constrained not only locally by the autoregressive loss $\mathcal{L}_{\text{AR}}$ from the next token prediction and LLM codebook matching, but also globally through visual alignment loss $\mathcal{L}_{\text{GVA}}$, which modulates its latent space using externally well-trained representations. We extract visual representations from a pretrained foundational model and further aggregate these features to obtain semantically enriched representations for alignment.
@@ 2.1 Visual Generation Models
Diffusions The impressive success of diffusion models [44, 42, 47, 30, 18] on image generation tasks has revolutionized the image generation paradigm. Recently, DiTs [35] and U-ViT [3] replace or integrate U-Net with Transformers, inspiring the development of models such as SD3 [15], Imagen3 [2], and Pixart [7], which achieve new state-of-the-art performance in image generation. However, diffusion models need to rely on separate language models (e.g., CLIP [38] or T5 [39]) as text encoder, and do not allow for uniform modeling of text and image data.
Autoregressive models Early pioneering works, VQ-VAE [53], VQ-GAN [14] and DALL-E [41], demonstrated the potential of AR models for image generation tasks. Subsequent works such as VQVAE-2 [43] and RQ-Transformer [21] also follow a raster-scan manner and enhance image generation performance through extra scales or stacked codes. Parti [60], based on the architecture of ViT-VQGAN [59], scales the autoregressive transformer to 20B parameters, demonstrating promising high-fidelity image generation results. Recent works achieve performance comparable to diffusion models by employing scale modeling [52] and eliminating vector quantization [25]. Additionally, masked prediction autoregressive models [5, 6, 26] employ BERT-like [11] masked prediction modeling, improving generation efficiency and quality.
@@ 2.2 LLMs for Text-to-Image Generation
LLM-based text-to-image generation models have attracted researchers’ attention, aiming to replicate their success in the field of text generation. Early work leverages diffusion models as tools to extend LLMs [12, 16, 49, 56]. These works utilize LLMs as feature extractors to guide additional diffusion generators for visual generation. Such models are complex to design and do not fully unleash the generative potential of the LLM for visual generation. Recent works [48, 50, 31, 32, 27] attempt to unify text and image modeling within a single LLM, enhancing generation performance through tokenizer optimization [48], early fusion modeling [50], and flexible resolution modeling [27]. These works discretize both text and images into tokens, which are then fed into the LLM for sequence modeling by next token prediction. However, the local constraints provided by next-token prediction struggle to bridge the significant cross-modal gap between language and images. Further, Transfusion [62] introduces bidirectional attention to model global structure via patch diffusion, while Show-O [58] employs masked token modeling to unify local and global dependencies. Although effective, such models typically deviate from the standard LLM framework. Unlike these approaches, our framework introduces global constraints through external visual representation alignment during training while preserving the powerful next-token prediction paradigm.
@@ 3 Proposed Method
We aim to achieve high-quality image generation without altering the next-token prediction paradigm of LLMs. We argue that the inherent cross-domain gaps in LLMs pose significant challenges for image generation tasks when the model lacks the ability to learn global features. To address this, we propose a Autoregressive Representation Alignment (ARRA) framework (see Fig. 2), which leverages the inherent representation capabilities of well-pretrained foundation models to facilitate the training of complex text-to-image autoregressive generation. Our framework is applied exclusively during training, without affecting inference, enabling the generation of high-quality images with exceptional semantic consistency in a cost-effective manner while accelerating convergence speed.
@@ 3.1 Overview
Our goal is to train an autoregressive model $\mathcal{M}_{\theta}$ leveraging representations derived from an external foundation visual encoder $\mathcal{E}_{F}$. $\mathcal{M}_{\theta}$ takes a text prompt $T$ as input and generates a target image $I$. During training, $T$ and $I$ are first tokenized into token sequences $s_{T}$ and $s_{I}$, respectively. These token sequences are then used to train the transformer-based autoregressive model $\mathcal{M}_{\theta}$. Meanwhile, foundation visual encoder $\mathcal{E}_{F}$ encodes $I$ into the global visual representation $f_{GF}$, which is used to align with the feature $f_{A}$ extracted from $x_{t}$ by $\mathcal{M}_{\theta}$. During image generation, the alignment module is removed, and image tokens are generated by $\mathcal{M}_{\theta}$ through next-token prediction. Finally, the output image tokens are decoded into pixel space by a image decoder to produce the target image $I$. We describe autoregressive modeling process in Section 3.2 and detail our autoregressive representation alignment framework in Section 3.3.
3.2 Autoregressive Modeling via Next-Token Prediction
The autoregressive architecture comprises two core components: (1) A transformer-based autoregressive model $\mathcal{M_{\theta}}$ for probabilistic modeling of token sequences. (2) A VQ-based model [14] with encoder $\mathcal{E}$, quantizer $\mathcal{Q}$ and decoder $\mathcal{D}$ for transformation between image pixels and discrete token sequences. Formulation involves following two parts:
@@ 3.2 Autoregressive Modeling via Next-Token Prediction
Next token prediction modeling. We combine text sequences $s_{T}$ and image sequences $s_{I}$ to obtain discrete tokens $x=\{x_{1},x_{2},x_{3},\dots,x_{n}\}$, where $x_{n}$ is an integer from a tokenizer’s vocabulary $V$. The next-token autoregressive posits the probability of current token $x_{t}$ depends only on its prefix $(x_{1},x_{2},x_{3},\dots,x_{t-1})$. The likelihood of sequence modeling can be expressed as:
@@
The autoregressive model $\mathcal{M_{\theta}}$ formulates the generative task as predicting the distribution of the next token and optimizes the likelihood $p(x)$ through cross-entropy (CE) loss:
@@
During training, the autoregressive model $\mathcal{M}_{\theta}$ relies on the previous tokens $x_{<t}$ to predict the next token $x_{t}$, where the hidden state of $x_{t}$ in the $i$-th layer of $\mathcal{M}_{\theta}$ is denoted by $f_{L}^{i}$. In the final layer, the hidden state $f_{L}^{-1}$ are then passed through the LLM head to compute the probability distribution $p_{\theta}$ for $x_{t}$. In the original autoregressive model, $p_{\theta}$ is constrained only by the local context of a single token (i.e. $x_{t}$), lacking the ability to capture global information. This limitation restricts the model’s capacity to capture complex cross-modal relationships. To address this problem, we propose Autoregressive Representation Alignment in Section 3.3.

### v4

@@ Abstract
We present Autoregressive Representation Alignment (ARRA), a new training framework that unlocks global-coherent text-to-image generation in autoregressive LLMs without architectural modifications. Different from prior works that require complex architectural redesigns, ARRA aligns LLM’s hidden states with visual representations from external visual foundational models via a global visual alignment loss and a hybrid token, <HYBNEXT>. This token enforces dual constraints: local next-token prediction and global semantic distillation, enabling LLMs to implicitly learn spatial and contextual coherence while retaining their original autoregressive paradigm. Extensive experiments validate ARRA’s plug-and-play versatility. When training T2I LLMs from scratch, ARRA reduces FID by 16.6% (ImageNet), 12.0% (LAION-COCO) for autoregressive LLMs like LlamaGen, without modifying original architecture and inference mechanism. For training from text-generation-only LLMs, ARRA reduces FID by 25.5% (MIMIC-CXR), 8.8% (DeepEyeNet) for advanced LLMs like Chameleon. For domain adaptation, ARRA aligns general-purpose LLMs with specialized models (e.g., BioMedCLIP), achieving an 18.6% FID reduction over direct fine-tuning on medical imaging (MIMIC-CXR). These results demonstrate that training objective redesign, rather than architectural modifications, can resolve cross-modal global coherence challenges. ARRA offers a complementary paradigm for advancing autoregressive models. The code is available at https://github.com/HKU-HealthAI/ARRA.
@@ 1 Introduction
Figure 1: ARRA enables high-quality text-to-image generation through a redefined training objective that promotes global coherence. (a)(c) Traditional next-token prediction (NTP)-based LLMs rely solely on the autoregressive loss (AR loss) of the next token <NEXT> for local constraints. (b)(d) ARRA constructs a hybrid token <HYBNEXT>, which is aligned by introducing external global visual representation, ensuring that <HYBNEXT> is constrained both locally by the AR loss and globally by the global visual alignment loss (GVA loss). (e)(f) ARRA demonstrates advantages in semantic consistency and visual continuity.
@@
Large language models (LLMs) (achiam2023gpt; yang2025qwen3; guo2025deepseek) and subsequent multimodal LLMs (MLLMs) (team2023gemini; wang2024emu3; liu2024improved; zhu2025internvl3) have revolutionized the field of generative AI. These models, built on the autoregressive (AR) paradigm, show remarkable scalability and generalization capabilities in complex understanding tasks through a simple yet powerful next-token prediction framework.
@@
Inspired by the success of LLMs, some researchers seek to replicate autoregressive “next-token prediction” paradigm for text-to-image generation. This paradigm is directly adopted in DALL·E (ramesh2021zero), Parti (yu2022scalingautoregressivemodelscontentrich), and LlamaGen (sun2024autoregressivemodelbeatsdiffusion) for image generation, where images are treated as sequences of discrete tokens. However, although “next-token prediction” paradigm excels in language tasks, where local dependencies naturally align with sequential structure, it struggles to bridge the significant cross-modal gap between language and images. As shown in Fig. 1(c), optimizing only for local next-token prediction forces the model to focus on isolated token-level features, neglecting the global coherence required for spatially structured visual content. This sometimes leads to fragmented parts on generated images, such as misaligned ribs in X-rays, where fine-grained details fail to harmonize into a unified whole. It can also cause semantic mismatches, as shown in Fig. 1(e), where global information is not maintained, leading to inconsistencies in the generated images.
@@
Recognizing this limitation, recent efforts aim to inject global constraints into autoregressive frameworks to fully unlock the potential of LLMs in image generation (zhou2024transfusion; xie2024show; tian2024visual). Bidirectional attention mechanisms are introduced in Transfusion (zhou2024transfusion) and Show-O (xie2024show) to model global image structure through patch diffusion and mask token modeling, respectively. These methods achieve promising results in generating high-quality images and demonstrate the potential of LLMs for multimodal generation. However, they rely on architectural modifications, such as cross-modal attention layers or grafted diffusion modules. While effective, such adaptations often deviate from standard LLM frameworks, limiting their compatibility with pretrained LLMs that excel under pure autoregressive paradigms. For instance, repurposing an off-the-shelf LLM for text-to-image generation would require retraining these modified components, losing benefits of existing scaling laws and generalization capabilities. This practical constraint raises a critical question: Can we unlock the full potential of LLMs for text-to-image generation without altering the original architecture or inference mechanism?
@@
We address this by proposing Autoregressive Representation Alignment (ARRA), a novel training framework that redefines how LLMs learn text-to-image generation. Different from prior works that modify architectures (e.g., adding attention layers or diffusion modules), ARRA preserves the original LLM framework while injecting global constraints directly into the training objective. Our key insight is simple: global coherence does not require architectural complexity; it can instead be achieved through a redefined training paradigm. Specifically, ARRA augments the standard autoregressive loss with a global visual alignment loss that aligns the LLM’s latent representations with semantic guidance from pretrained foundational models (Fig. 2). To bridge local and global learning, we introduce a hybrid token, <HYBNEXT>, which serves as a bidirectional anchor. Locally, it predicts the next token via standard codebook indices. Globally, its latent embedding aligns with compressed visual features extracted from external models (e.g., BioMedCLIP (zhang2023biomedclip) or MedSAM (MedSAM)) via our novel global visual alignment loss. By distilling rich semantic features (e.g., spatial relationships, object coherence) from external models into the <HYBNEXT> token during training, ARRA enables autoregressive sequences to implicitly learn global structure. More crucially, this alignment occurs only during training, leaving the LLM’s inference process untouched and preserving its inference-time efficiency.
Our experiments demonstrate the versatility of ARRA framework for both natural and medical image generation tasks, achieving improvements without architectural modifications. ARRA supports three key capabilities on AR LLMs: (1) ARRA enhances training T2I LLMs from scratch. Applied to LlamaGen (sun2024autoregressivemodelbeatsdiffusion) with varying parameter scales, it consistently improves generation performance and exhibits strong scalability. (2) ARRA effectively transforms pretrained text-generation-only LLMs into T2I generators. When integrated into LLMs without image generation capabilities, such as Chameleon (team2024chameleon), ARRA yields steady improvements. (3) ARRA facilitates adaptation of general-purpose generative models to special domains. By integrating domain-specific priors (e.g., BioMedCLIP, MedSAM) into LLMs with image generation capabilities, such as Lumina-mGPT (liu2024lumina), ARRA substantially outperforms direct fine-tuning. These capabilities confirm the plug-and-play flexibility of ARRA framework.
@@
(i) We propose Autoregressive Representation Alignment, a novel training framework that redefines how LLMs learn text-to-image generation by decoupling global structure learning from model design. By aligning training objective with external representations, ARRA resolves local dependency limitations in LLMs while retaining original architectures and inference efficiency.
(ii) We introduce the <HYBNEXT> token, a novel mechanism that bridges local next-token prediction with global semantic alignment via distillation from external models (e.g., BioMedCLIP or MedSAM), enabling implicit learning of spatial and contextual relationships.
(iii) We provide a detailed experimental analysis, offering comprehensive insights into the selection of alignment tokens, aggregation strategy, and external representations. These findings serve as a practical guide for the effective utilization of representation alignment.
(iv) ARRA offers plug-and-play flexibility, facilitating training T2I LLMs from scratch, transforming pretrained text-generation-only LLMs into T2I generators, and adapting general generative models to specific domains, all without architectural modifications. These capabilities are validated on both natural and medical image generation tasks using advanced AR models.
Figure 2: Proposed ARRA Framework. We define the next token predicted in the autoregressive sequence as the “HYBRID next token”, denoted as <HYBNEXT>. During training, <HYBNEXT> is constrained not only locally by the autoregressive loss $\mathcal{L}_{\text{AR}}$ from the next token prediction and LLM codebook matching, but also globally through visual alignment loss $\mathcal{L}_{\text{GVA}}$, which modulates its hidden states using externally well-trained representations. We extract visual representations from a pretrained foundational model and further aggregate these features to obtain semantically enriched representations for alignment.
@@ 2.1 Visual Generation Models
Diffusions. The success of diffusion models has revolutionized the image generation paradigm (qu2015pixel; qu2019wavelet; rombach2022SD; liu2024residual; xie2025dvg). DiTs (peebles2023scalable) show strong scalability by replacing or integrating U-Net with Transformers. This inspires subsequent models such as SD3 (esser2024SD3), Imagen3 (baldridge2024imagen), which achieve new state-of-the-art performance in image generation. Recent work REPA (yu2025repa) explores enhancing diffusion models with external representations. It utilizes patch-wise representations alignment, matching each patch-level hidden state of the diffusion transformer with corresponding patch tokens from an external encoder. This alignment improves generation performance without modifying architecture. However, REPA’s strategy is not directly compatible with AR models. AR models generate image tokens sequentially and do not produce all patch tokens simultaneously during training, making patch-wise alignment infeasible. Different from this, ARRA introduces a novel hybrid token to bridge local next-token prediction and global alignment, enabling effective integration of the alignment mechanism into AR architectures. For further discussion, please refer to Appendix 2.
Autoregressive models. Early pioneering works, VQ-VAE (van2017neural), VQ-GAN (esser2021taming) and DALL-E (ramesh2021zero), demonstrated the potential of AR models for image generation. Subsequent works such as RQ-Transformer (lee2022autoregressive) also follows a raster-scan manner and enhances image generation performance through extra scales or stacked codes. Recent works achieve performance comparable to diffusion models by employing scale modeling (tian2024visual) and eliminating vector quantization (li2024autoregressive; fan2024fluid). Additionally, masked prediction autoregressive models (chang2022maskgit; li2024mar) employ BERT-like (devlin2019bert) masked prediction modeling, improving generation efficiency and quality.
@@ 2.2 LLMs for Text-to-Image Generation
Currently, researchers pay attention to LLM-based text-to-image generation models, aiming to replicate the success of LLMs in language tasks. Early works leverage diffusion models as tools to extend LLMs (dong2024dreamllm; ge2024seed; sun2024generative). These works utilize LLMs as feature extractors to guide diffusion generators for visual generation. Such models are complex to design and do not fully unleash the generative potential of the LLM for visual generation. Recent works (sun2024autoregressivemodelbeatsdiffusion; team2024chameleon; lu2024unified; liu2024lumina) attempt to unify text and image modeling within a single LLM, enhancing generation performance through tokenizer optimization (sun2024autoregressivemodelbeatsdiffusion), early fusion modeling (team2024chameleon), and flexible resolution modeling (liu2024lumina). These works discretize both text and images into tokens, which are then fed into the LLM for sequence modeling by next token prediction. However, the local constraints provided by next-token prediction struggle to bridge the cross-modal gap between language and images.
@@ 3 Proposed Method
We aim to achieve high-quality image generation without altering the next-token prediction paradigm of LLMs. We argue that the inherent cross-domain gaps in LLMs pose significant challenges for image generation tasks when the model lacks the ability to learn global features. To address this, we propose an Autoregressive Representation Alignment (ARRA) framework (see Fig. 2), which leverages the inherent representation capabilities of well-pretrained foundation models to facilitate the training of complex text-to-image autoregressive generation. Our framework is applied only during training, without affecting inference. It can enable the generation of high-quality images with exceptional semantic consistency in a cost-effective manner.
@@ 3.1 Overview
Our goal is to train an autoregressive model $\mathcal{M}_{\theta}$ leveraging representations derived from an external foundation visual encoder $\mathcal{E}_{F}$. $\mathcal{M}_{\theta}$ takes a text prompt $T$ as input and generates a target image $I$. During training, $T$ and $I$ are first tokenized into token sequences $s_{T}$ and $s_{I}$, respectively. These token sequences are then used to train the transformer-based autoregressive model $\mathcal{M}_{\theta}$. Meanwhile, foundation visual encoder $\mathcal{E}_{F}$ encodes $I$ into the global visual representation $f_{GF}$, which is used to align with the feature $f_{A}$ extracted from $x_{t}$ by $\mathcal{M}_{\theta}$. During image generation, the alignment module is removed, and image tokens are generated by $\mathcal{M}_{\theta}$ through next-token prediction. Finally, the output image tokens are decoded into pixel space by an image decoder to produce the target image $I$. We describe the autoregressive modeling process in Section 3.2 and detail our autoregressive representation alignment framework in Section 3.3.
3.2 Modeling via Next-Token Prediction
The autoregressive architecture comprises two core components: (1) A transformer-based autoregressive model $\mathcal{M_{\theta}}$ for probabilistic modeling of token sequences. (2) A VQ-based model (esser2021taming) with encoder $\mathcal{E}$, quantizer $\mathcal{Q}$ and decoder $\mathcal{D}$ for transformation between image pixels and discrete token sequences. Formulation involves the following two parts:
@@ 3.2 Modeling via Next-Token Prediction
Next token prediction modeling. We combine text sequences $s_{T}$ and image sequences $s_{I}$ to obtain discrete tokens $x=\{x_{1},x_{2},x_{3},\dots,x_{n}\}$, where $x_{n}$ is an integer from a tokenizer’s vocabulary $V$. The next-token prediction paradigm posits the probability of current token $x_{t}$ depends only on its prefix $(x_{1},x_{2},x_{3},\dots,x_{t-1})$. The likelihood of sequence modeling can be expressed as:
@@
The autoregressive model $\mathcal{M_{\theta}}$ formulates the generative task as predicting the distribution of the next token and optimizes the likelihood $p_{\theta}(x)$ through cross-entropy (CE) loss:
@@
During training, the autoregressive model $\mathcal{M}_{\theta}$ relies on the previous tokens $x_{<t}$ to predict the next token $x_{t}$, where the hidden state of $x_{t}$ in the $i$-th layer of $\mathcal{M}_{\theta}$ is denoted by $f_{L}^{i}$. In the last layer, the hidden state $f_{L}^{-1}$ is then passed through the LLM head to compute the probability distribution $p_{\theta}$ for $x_{t}$. In the original autoregressive model, $p_{\theta}$ is constrained only by the local context of a single token (i.e., $x_{t}$), lacking the ability to capture global information. This limitation restricts model’s capacity to capture complex cross-modal relationships. To address this problem, we propose Autoregressive Representation Alignment.